OpenStudy (anonymous):

If t varies as v, and t = 2 4/7 when v = 13/14 , find v when t = 2 1/4. v= Direct variation

OpenStudy (dumbcow):

direct variation: v = kt, where k is a constant plug in the given v and t to solve for k 13/14 = 18/7 k k = 13/36 --> v = 13/36 t now plug in any t to find the resulting v

OpenStudy (anonymous):

do I multiply 18/7 by 13/36?

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

gawd I feel like I'm insulting you...

OpenStudy (anonymous):

use the unitary method t v 24/7 13/14 21/4 hence v={(13/14)/(24/7)}*(21/4) v=19/64

OpenStudy (anonymous):

thank you Sarkar!:)

OpenStudy (anonymous):

yw